# VisionTS: Visual Masked Autoencoders Are Free-Lunch Zero-Shot Time Series Forecasters

### v3

@@ 6 Conclusion
This paper presents work whose goal is to advance the field of time series forecasting. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@
Appendix A Extended Related Work
Text-based TSF Foundation Models
Appendix B Details of Experiments
B.1 Benchmark and baselines
@@ B.2 Periodicity selection
Appendix C Zero-Shot Forecasting
C.1 Hyperparameters
@@
C.2 Full forecasting results of the long-term TSF benchmark
@@
C.3 Comparison of TimesFM and LLMTime
@@
C.4 Comparison of traditional methods
@@
C.5 Full forecasting results of the Monash TSF benchmark
@@ C.6 Impact of backbones
C.7 Impact of the different image encoding strategies
@@ C.8 Hyperparameter analysis
Appendix D Full-Shot Forecasting
D.1 Training details
@@
D.2 Full results and standard deviations
@@
D.3 Ablation study and fine-tuning strategy comparison

### v4

@@ 6 Conclusion
Acknowledgments
Research work mentioned in this paper is supported by State Street Zhejiang University Technology Center. We would also like to thank reviewers for their valuable comments.
@@ Acknowledgments
This paper presents work whose goal is to advance the field of time series forecasting. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@
Appendix A Details of Experiments
A.1 Benchmark and baselines
@@ A.2 Periodicity selection
Appendix B Zero-Shot Forecasting
B.1 Hyperparameters
@@
B.2 Full forecasting results of the long-term TSF benchmark
@@
B.3 Comparison of TimesFM and LLMTime
@@
B.4 Comparison of traditional methods
@@
B.5 Comparison of concurrent works
Table 13: Comparison of Time-MoE and TTM in the zero-shot setting. We report the base and large model results for Time-MoE, as the ultra model weights are not yet released. For TTM, we used the official HuggingFace model for replication. The following table summarizes the performance of various zero-shot foundation models.
VisionTS Time-MoE (base) Time-MoE (large) TTM (v1)
ETTh1 MSE 0.390 0.400 0.394 0.398
MAE 0.414 0.424 0.419 0.421
ETTh2 MSE 0.333 0.366 0.405 0.348
MAE 0.375 0.404 0.415 0.393
ETTm1 MSE 0.374 0.394 0.376 0.520
MAE 0.372 0.415 0.405 0.479
ETTm2 MSE 0.282 0.317 0.316 0.312
MAE 0.321 0.365 0.361 0.348
Electricity MSE 0.207 (data leakage) (data leakage) 0.201
MAE 0.294 (data leakage) (data leakage) 0.293
Weather MSE 0.269 0.265 0.270 0.234
MAE 0.292 0.297 0.300 0.266
Average MSE 0.309 - - 0.335
MAE 0.345 - - 0.367
[…]
Table 14: Comparison of CALF in both zero-shot and full-shot settings.
VisionTS (zero-shot) VisionTS (full-shot) CALF
ETTh1 MSE 0.390 0.395 0.432
MAE 0.414 0.409 0.428
ETTh2 MSE 0.333 0.336 0.349
MAE 0.375 0.382 0.382
ETTm1 MSE 0.374 0.338 0.395
MAE 0.372 0.367 0.390
ETTm2 MSE 0.282 0.261 0.281
MAE 0.321 0.319 0.321
Electricity MSE 0.207 0.156 0.175
MAE 0.294 0.249 0.265
Weather MSE 0.269 0.227 0.250
MAE 0.292 0.262 0.274
Average MSE 0.309 0.286 0.314
MAE 0.345 0.331 0.343
We compare our work with other concurrent TSF methods. Table 13 presents the comparison from Time-MoE (Shi et al., 2024) and TTM (Ekambaram et al., 2024), and Table 14 shows the comparison with CALF (Liu et al., 2025), which is the existing SOTA LLMs-based time series forecasting work. These findings highlight the promising potential of vision models in TSF scenarios.
B.6 Full forecasting results of the Monash TSF benchmark
@@ B.7 Impact of backbones
B.8 Impact of the different image encoding strategies
@@ B.9 Hyperparameter analysis
Appendix C Full-Shot Forecasting
C.1 Training details
@@
C.2 Full results and standard deviations
@@
C.3 Ablation study and fine-tuning strategy comparison